OpenStudy (anonymous):

is (0,3) a solution to the equation y=x3?

OpenStudy (nincompoop):

is it?

OpenStudy (anonymous):

Please ask this question in the Math section

OpenStudy (nincompoop):

\[(x, y) \rightarrow (0, 3): y = x^3 \rightarrow 3 = 0^3 \] does that look like the right solution to your problem?

OpenStudy (nincompoop):

does zero raised to the third power equal 3?

OpenStudy (nincompoop):

that is what your problem is essentially asking you need to be able to interpret these things better by understanding the concepts

OpenStudy (anonymous):

0,3) means x = 0 and y = 3 y=x + 3 y=0 + 3 ... replace x with 0 (since x = 0) 3 = 0 + 3 ... replace y with 3 (since y = 3) 3 = 3 since that's true, this means that (0,3) is a solution to y = x+3 put it in the math section next time.

OpenStudy (nincompoop):

@daysofashley y = x^3 not y = x+3

OpenStudy (anonymous):

@nincompoop thanks for that.